# Tailoring Personality Traits in Large Language Models via Unsupervisedly-Built Personalized Lexicons

### v2

@@ Abstract
Personality plays a pivotal role in shaping human expression patterns, thus regulating the personality of large language models (LLMs) holds significant potential in enhancing the user experience of LLMs. Previous methods either relied on fine-tuning LLMs on specific corpora or necessitated manually crafted prompts to elicit specific personalities from LLMs. However, the former approach is inefficient and costly, while the latter cannot precisely manipulate personality traits at a fine-grained level. To address the above challenges, we have employed a novel Unsupervisedly-Built Personalized Lexicons (UBPL) in a pluggable manner during the decoding phase of LLMs to manipulate their personality traits. UBPL is a lexicon built through an unsupervised approach from a situational judgment test dataset (SJTs4LLM). Users can utilize UBPL to adjust the probability vectors of predicted words in the decoding phase of LLMs, thus influencing the personality expression of LLMs. Extensive experimentation demonstrates the remarkable effectiveness and pluggability of our method for fine-grained manipulation of LLM’s personality.
Tailoring Personality Traits in Large Language Models via Unsupervisedly-Built Personalized Lexicons
Tianlong Lithanks:   These authors contributed equally., Shihan Dou11footnotemark: 1, Changze Lv, Wenhao Liu Jianhan Xu, Muling Wu, Zixuan Ling, Xiaoqing Zhengthanks:   Corresponding author., Xuanjing Huang School of Computer Science, Fudan University, Shanghai, China {tlli22,shdou21,czlv22,whliu22}@m.fudan.edu.cn {zhengxq,xjhuang}@fudan.edu.cn
@@ 1 Introduction
Figure 1: The radar chart illustrates the effectiveness of the UBPL in manipulating the personality traits of the Llama2-13b-chat. In the chart, the red dashed line denotes the expressive levels of five personality traits without UBPL. The light orange area indicates the range within which the model’s personality traits can be manipulated in our main experimental setting ($\alpha=1$, $|\beta_{t}|\leq 1$ and $\beta_{\neq t}=0$, where $\alpha$, $\beta$ are user-controlled hyperparameters that adjust the effects of UBPL). When users change the hyperparameters, the range of control is not limited to the light orange area shown in the chart.
@@
With rapid expansion in scale, LLMs demonstrate superior capabilities for high-quality text generation and revolutionize traditional natural language processing tasks (Wei et al., 2022). This forefront development has sparked concerns about the security, ethics, and potential hallucinatory issues associated with the proliferation of AI-generated content (AIGC), while also fueling a substantial rise in user demand for personalized agent services based on LLMs (Hagendorff, 2023). Personalized agent models can tailor their expression of personality patterns based on user preferences, making it closely aligned with user habits and thereby enhancing the overall user experience. This is achieved by regulating the style and behavior patterns of their interactions with users, which are often referred to as the "personality" of LLMs (Allport, 1961). Prior studies have also defined this personality as the presence of stable and internally consistent patterns of behavior in LLMs and found that different LLMs have different personalities (Miotto et al., 2022; Caron and Srivastava, 2022; Karra et al., 2022).
@@
Presently, there are two main effective methods to alter the personality of LLMs: fine-tuning and prompt engineering. While the former (Karra et al., 2022) can effectively change the personality of LLMs in specific dimensions, it is not only inefficient (requiring resource-consuming parameter updates for each model) but also incapable of achieving finer-grained control. The latter, while not requiring adjustments to model parameters, still falls short in achieving fine-grained control over the personality of LLMs (Jiang et al., 2022; Safdari et al., 2023; Pan and Zeng, 2023; Tu et al., 2023).
@@
To address the problem of these above methods, we propose to leverage an Unsupervisedly-Built Personalized Lexicon (UBPL) to intervene in the decoding phrase of LLMs in a pluggable manner, enabling fine-grained control over their personality. Figure 1 shows the remarkable effect of our method to manipulate the personality of LLMs. UBPL is a lexicon built through an unsupervised approach from a Situational Judgment Tests (SJTs) dataset (SJTs4LLM) we constructed. SJTs4LLM is built based on the Big Five personality theory (De Raad, 2000), which consists of a question set and an answer set (McDaniel et al., 2007). The question set is generated by GPT-4 and manually filtered to assess the levels of the different personalities of LLMs. Following previous studies (Karra et al., 2022; Caron and Srivastava, 2022), in the process of assessing the personalities of LLMs, the models’ responses to SJTs4LLM questions were recorded and subjected to statistical analysis by using a five-dimensional Likert scale. The answer set contains the answer text with different personality traits and is used to build a UBPL dedicated to an LLM.
@@
Our method not only eliminates the need for resource-intensive full-parameter fine-tuning of LLMs but also allows users to adjust only a few parameters to achieve fine-grained manipulation of the different personalities of LLMs. We have done extensive experiments using six popular LLMs to demonstrate the pluggable convenience and remarkable effectiveness of our method. Our contribution can be summarized as follows:
@@
  • We propose a novel method for exerting control over the personality of Language Models, leveraging UBPL to intervene at the decoding phase. This method enables fine-grained controllability over the behavior of LLMs without necessitating updates to the model parameters.
@@
  • We constructed a new dataset inspired by the concept of Situational Judgment Tests, marking the pioneering effort in datasets especially created for the evaluation of LLM’s personality. Diverging from conventional direct psychological questionnaires, this indirect approach shows enhanced intuitiveness and reliability in the assessment of personality traits.
@@
  • Extensive experiments were conducted with different LLMs on the SJTs4LLM dataset, revealing that the UBPL method demonstrates notable effectiveness in achieving both enhanced efficiency and finer-grained control over LLM’s personality traits.
@@ 2.1 The Big Five
In the realm of research within the field of psychological measurement, various classification systems for human personality traits exist, such as the Sixteen Personality Factors (16PF) (Cattell and Mead, 2008) and Myers–Briggs Type Indicator (MBTI) (Miles and Hempel, 2004). Among them, the Big Five (De Raad, 2000) stands out as a widely embraced model for personality trait modeling, effectively defining and describing the inherent behavioral patterns within individuals. This theory quantifies human personality traits into five dimensions: Openness(OPE), Conscientiousness(CON), Extraversion(EXT), Agreeableness(AGR), and Neuroticism(NEU). For a detailed description of each personality trait and how they relate to each other, please refer to Appendix A.
@@ 2.2 Methods for controlling LLMs personality
Prompt engineering. Jiang et al. (2022) proposed the method of Personality Prompting ($P^{2}$) to construct the prompts that can effectively induce a specific personality through multiple steps; Safdari et al. (2023) utilized a novel prompting methodology grounded in lexical hypotheses (Goldberg, 1981) to effectively shape personalities in LLMs, encompassing both single-trait and multi-trait dimensions. In addition, Pan and Zeng (2023) and Tu et al. (2023) also attempted to change the personality of LLMs through prompt engineering.
@@ 2.3 Situation Judgment Tests
Situation Judgment Tests (SJTs) have been described as "psychometric alchemy" and are typically viewed as contextual selection procedures that assess a candidate’s responses to various relevant work situations, serving as a predictive tool (Lievens and Motowidlo, 2016; Bledow and Frese, 2009). The advantage of SJTs is that their validity and incremental validity are higher than those of cognitive ability and personality tests because SJTs do not require the subject to give a direct answer to the question, but give the subject a situational premise to evaluate a certain characteristic of the subject through the side of the subject’s choice (such as the Q: Your partner suggests creating a YouTube channel to document and share your unique hobbies or interests. Are you willing to share your passion with a wider audience?) (Lievens et al., 2008). Compared to the direct questionnaire tests used in previous jobs (such as the Q: Are you a risk-taker and unconventional person?) This feature of SJTs can effectively bypass the preference defenses of LLMs, resulting in more trustworthy personality assessments (Figure 2 shows another example of SJTs).
Figure 2: Direct questionnaires vs. Situational Judgment Tests (SJTs). The questions in the direct questionnaires are often abstract, making it challenging for models trained through Reinforcement Learning from Human Feedback (RLHF) and instruction alignment to generate the desired responses. In contrast to direct questionnaires, SJTs present a unique approach by adopting a "role-playing" hypothetical perspective to deceive and induce the model’s responses. Subsequently, we can indirectly assess the extent to which the model manifests personality traits based on these responses.
@@ 3 Method
The proposed UBPL in our study can be integrated into any open-source LLM in a pluggable manner and effectively adjust the diverse personality traits that LLMs exhibit (based on the Big Five theory). Figure 3 shows the specifics of our method.
The first step is to build UBPL in an unsupervised manner, using SJTs4LLM, the first SJTs dataset we built to assess the personality of LLMs. Employing the tokenizer of a chosen model, such as Llama2 for illustration, we tokenize each text in the answer set of SJTs4LLM. Subsequently, we conduct a subword-level statistical analysis of the tokenized texts, categorizing the obtained subwords based on the personality trait themes to which each text belongs. Following this, we assign values within UBPL to each subword (also serving as the "key" of UBPL) corresponding to the respective personality lists. After processing all texts in the answer set, we normalize and scale the values within UBPL based on distinct personality themes.
@@
In the second step, we use UBPL to manipulate personality in the decoding phase of the LLMs. By default, the LLMs adopt the Top-p nucleus sampling strategy. During LLMs’ decoding, we concatenate additional personalized probability combinations from UBPL after filtering out low-probability predicted subwords with cumulative probabilities below a threshold $p$. Subsequently, normalization and multinomial sampling procedures are applied.
@@
Figure 3: Illustration of our methods. Constructing SJTs4LLMs: We employed Prompt1 (P1) to prompt GPT-4 for generating responses, which were subsequently curated through manual screening to construct the question set for SJT4LLMs. Subsequently, models such as GPT-4 were engaged in a "role-playing" manner using Prompt2 (P2) to generate answers with diverse personality traits tailored to this question set, thus forming the answer set for SJT4LLMs; Building UBPL Unsupervisedly: Initially, we employ the tokenizer of LLMs (sp(·)) to tokenize each answer text in the answer set of SJTs4LLMs. Subsequently, we categorize the obtained sub-words based on the personality trait theme to which the answer belongs. Within each personality trait category, cool-toned words signify a low expression level of that trait, while warm-toned words indicate a high expression level. On this basis, UBPL is built through two steps: assignment, normalization & scaling; Decoding with Personality Injection: We employ Prompt3 (P3) to prompt the model to answer the question set of SJTs4LLMs. During the model’s decoding process utilizing Top-P nucleus sampling, we used UBPL each time to change the probability vector of the next prediction word and finally changed the expression degree of personality traits of the model’s answers.
3.1 Building UBPL Unsupervisedly
@@
When building UBPL, we utilized the answer set of SJTs4LLM, which comprises subsets of personality trait answer texts with 10 distinct polarities (each of the 5 personality traits further refined into high and low subtraits).
@@
We denote this answer set as $A$, the tokenizer of the model as $sp(\cdot)$, the vocabulary of the model as $V$, and UBPL as $L$.
$L=(L_{key},L_{val})$ (1)
@@
where $L_{key}$ is initialized using $V$, and the initial values of $L_{val}$ are zero lists of length 5.
The entire construction process is divided into two parts: assignment operations and normalization and scaling operations.
In the first step, we tokenize the texts $A_{ij}$ in $A$:
$\text{sp}(A_{ij})=\{w_{1},w_{2},\ldots\}$ (2)
@@
Then, we define the personality trait index $t$: $t=\frac{i}{2}$. Finally, we perform the assignment on $L_{val}$:
$L^{t}_{val}=L^{t}_{val}+\begin{cases}+1&\text{if }i\%2=0\\ -1&\text{else}\end{cases}$ (3)
@@
$L^{t}_{val}=\{v_{1},v_{2},\ldots,v_{m},v_{m+1},\ldots,v_{m+n}\},$ (4)
where $m+n=|V|$, and for $k\leq m$, $v_{k}\geq 0$. When the personality trait is $t$, the averages of the positive value set and the negative value set of $L^{t}_{\text{val}}$ can be expressed as follows:
$\text{Avg}^{+}(t)=\frac{1}{m}\sum_{k=1}^{|V|}\max(0,L^{t}_{val,k})$ (5)
$\text{Avg}^{-}(t)=\frac{1}{n}\sum_{k=1}^{|V|}\min(0,L^{t}_{val,k})$ (6)
The second step involves normalizing and scaling the values in $L$. We define hyperparameters $S$, $M$, and $\alpha$ to control the degree of normalization and scaling. The $N$ in Equation 8 is employed to govern the normalization process, and its value is determined through a binary search. The termination condition is defined as the point at which the involvement of $N$ in the normalization of $L$ satisfies the condition specified in Inequality 9. The normalization and scaling process can be represented by the following mapping function:
$L^{t}_{val}\xrightarrow{(N_{t},S_{t})}F(L^{t}_{val},N_{t},S_{t})$ (7)
Specifically, it is expressed as:
@@
Where the value of $N$ is obtained through binary search to satisfy the following necessary conditions:
$\max\{|Avg^{-}(t)-M_{t}|,|Avg^{+}(t)-M_{t}|\}\leq\alpha$ (9)

### v3

# UPLex: Fine-Grained Personality Control in Large Language Models via Unsupervised Lexical Modulation

@@ Abstract
Personality is a crucial factor that shapes human communication patterns, thereby regulating the personalities of large language models (LLMs) holds significant potential in enhancing their user experiences. Previous approaches either relied on fine-tuning LLMs on specific corpora or required manually crafted prompts to evoke specific personalities from LLMs. However, the former is inefficient and costly, while the latter cannot precisely manipulate personality traits at a fine-grained level. To address these challenges, we propose UPLex, a method that uses an Unsupervisedly-Built Personalized Lexicon (UPL) during the decoding phase to manipulate LLM’s personality traits. UPL can be constructed from a newly built situational judgment test dataset in an unsupervised fashion, and used to modulate the personality expression of LLMs by dynamically altering their predicted probability of upcoming words in a pluggable fashion. Extensive experimentation demonstrates the remarkable effectiveness and pluggability of our method for fine-grained manipulation of LLMs’ personalities.
UPLex: Fine-Grained Personality Control in Large Language Models via Unsupervised Lexical Modulation
Tianlong Li, Wenhao Liu, Muling Wu, Shihan Dou, Zhenghua Wang, Changze Lv, Xiaohua Wang, Xiaoqing Zhengthanks:   Corresponding author., Xuanjing Huang School of Computer Science, Fudan University, Shanghai, China {tlli22,whliu22,mlwu22}@m.fudan.edu.cn {zhengxq,xjhuang}@fudan.edu.cn
Figure 1: The red dashed line denotes the expressive levels of the five personality traits of Llama2-13b-chat without UPL. At the same time, the green area represents the adjustable range of these traits with UPL in the main experiment. The personality names and their specific adjustable ranges are labeled. Additionally, we present a case in which the Openness trait is decreased using UPL.
@@ 1 Introduction
With rapid expansion in scale, LLMs demonstrate superior capabilities for high-quality text generation and revolutionize traditional natural language processing tasks (Wei et al., 2022; Kojima et al., 2022; Wei et al., 2021). This forefront development has sparked concerns about the safety, ethics, and potential hallucinatory issues associated with the proliferation of AI-generated content (AIGC), while also fueling a substantial rise in user demand for personalized agent services based on LLMs (Hagendorff, 2023; Zhang et al., 2023; Li et al., 2024). Personalized agent models can tailor their expression of personality traits according to user preferences, thereby closely aligning with user habits and enhancing the overall user experience. This is accomplished by modulating the interaction styles and behavior patterns, commonly called the “personalities” of LLMs (Allport, 1961; Jiang et al., 2023; Mao et al., 2023; Wang et al., 2023b). Prior studies have also defined this personality as the presence of stable and internally consistent patterns of behavior in LLMs and found that different LLMs have different personalities (Miotto et al., 2022; Caron and Srivastava, 2022; Karra et al., 2022; Bodroza et al., 2023).
@@
Recently, there have been two main effective methods to alter the personality of LLMs: fine-tuning and prompt engineering. While the former (Karra et al., 2022) can effectively modify the personality of LLMs in specific dimensions, it is not only inefficient (requiring resource-consuming parameter updates for each model) but also incapable of achieving finer-grained control. The latter, while not requiring adjustments to model parameters, still falls short in achieving fine-grained control over the personality of LLMs (Jiang et al., 2022; Safdari et al., 2023; Pan and Zeng, 2023; Tu et al., 2023).
To address the limitations of the above methods, we propose leveraging an Unsupervisedly-Built Personalized Lexicons (UPL) to intervene in the decoding phase of LLMs in a pluggable manner, enabling fine-grained control over their personalities. Figure 1 demonstrates the remarkable effectiveness of our method in manipulating LLM personalities.
@@
The UPL is constructed using an unsupervised approach from a Situational Judgment Tests (SJT) dataset (STD), based on the Big Five personality theory (De Raad, 2000). STD comprises a question set and an answer set (McDaniel et al., 2007). The question set is generated by GPT-4 and manually filtered to assess the levels of different personalities in LLMs. Following previous studies (Karra et al., 2022; Caron and Srivastava, 2022), in the process of assessing the personalities of LLMs, the models’ responses to STD questions were recorded and subjected to statistical analysis by using a five-dimensional Likert scale. The answer set contains texts with different personality traits and is used to build an UPL dedicated to an LLM.
@@
Our method not only obviates the need for resource-intensive fine-tuning of LLMs but also enables users to adjust a few parameters for fine-grained manipulation of LLMs’ various personalities. We have conducted extensive experiments with six popular LLMs to demonstrate the method’s pluggable convenience and remarkable effectiveness. The contribution of this study can be summarized as follows:
@@
  • We propose a novel method for exerting control over the personalities of LLMs, leveraging UPL to intervene at the decoding phase. This method enables fine-grained controllability over the personalities of LLMs without necessitating updates to the model parameters.
@@
  • We constructed a new dataset inspired by the concept of Situational Judgment Tests, marking the pioneering effort in datasets especially created for the evaluation of LLMs’ personalities. Diverging from conventional direct psychological questionnaires, this indirect approach shows enhanced intuitiveness and reliability in the assessment of personality traits.
@@
  • Extensive experiments were conducted with various LLMs on the STD, revealing that our method is notably effective in achieving enhanced efficiency and finer-grained control over the personality traits of LLMs.
@@ 2.1 The Big Five
In the field of psychometric research, there are various classification systems of personality traits, such as the Sixteen Personality Factors (16PF) (Cattell and Mead, 2008) and Myers–Briggs Type Indicator (MBTI) (Miles and Hempel, 2004). Among them, the Big Five (De Raad, 2000) stands out as a widely embraced model for personality trait modeling, effectively defining and describing the inherent behavioral patterns within individuals. This theory quantifies human personality traits into five dimensions: Openness(OPE), Conscientiousness(CON), Extraversion(EXT), Agreeableness(AGR), and Neuroticism(NEU). For a detailed description of each personality trait and how they relate to each other, please refer to Appendix B.
@@ 2.2 Methods for controlling LLMs personality
Prompt engineering. Jiang et al. (2022) proposed the method of Personality Prompting ($P^{2}$) to construct the prompts that can effectively induce a specific personality through multiple steps; Safdari et al. (2023) utilized a novel prompting methodology grounded in lexical hypotheses Goldberg (1981) to effectively shape personalities in LLMs, encompassing both single-trait and multi-trait dimensions. In addition, Pan and Zeng (2023) and Tu et al. (2023) also attempted to change the personality of LLMs through prompt engineering.
@@ 2.3 Situation Judgment Tests
Situation Judgment Tests (SJT) have been described as “psychometric alchemy” and are typically viewed as contextual selection procedures that assess a candidate’s responses to various relevant work situations, serving as a predictive tool (Lievens and Motowidlo, 2016; Bledow and Frese, 2009). SJT offer the advantage of having higher validity and incremental validity compared to cognitive ability and personality tests. This is because SJT do not ask subjects to provide direct answers. Instead, they present situational premises, allowing the evaluation of certain personality traits of the subjects based on their choices (such as the Q: Your partner suggests creating a YouTube channel to document and share your unique hobbies or interests. Are you willing to share your passion with a wider audience?) (Lievens et al., 2008). Compared to the direct questionnaire tests used in previous works (such as the Q: Are you a risk-taker and unconventional person?), this feature of SJT can effectively bypass the preference defenses of LLMs, resulting in more trustworthy personality assessments (Figure 2 shows another example of SJT).
Figure 2: Direct questionnaires vs. Situational Judgment Tests (SJT). The questions in the direct questionnaires are often abstract, making it challenging for models trained through Reinforcement Learning from Human Feedback (RLHF) and instruction alignment to generate the desired responses. In contrast to direct questionnaires, SJT present a unique approach by adopting a “role-playing” hypothetical perspective to deceive and induce the model’s responses. Subsequently, we can indirectly assess the extent to which the model manifests personality traits based on these responses.
@@ 3 Method
In this section, we elaborate on the method for constructing UPL and manipulating the diverse personality traits that LLMs exhibit through UPL. Refer to Figure 3 for an enhanced understanding of this section.
To build UPL in an unsupervised manner, we have constructed STD, an SJT dataset for assessing the personality of LLMs based on the Big Five theory. The UPL is formally a dictionary, where the keys are derived from the sub-words of a target model, and the values are obtained through word frequency analysis. We first tokenize each text in the answer set of STD with the tokenizer of the target model; Subsequently, we determined the initial value in the UPL based on the frequency of sub-words appearing in answer sets with different personality trait polarities. After processing all texts in the answer sets, we normalize and scale the values within UPL.
@@
When manipulating the personality of a target model with UPL, we adopt the Top-p nucleus sampling strategy to balance the original generation of LLMs and the impact of injected personality, which is consistent with the setting of Factual-Nucleus Sampling Lee et al. (2022). During the LLMs’ decoding process, we concatenate additional personalized probability distributions from the UPL after filtering out low-probability predicted subwords with cumulative probabilities below a threshold. Subsequently, normalization and multinomial sampling procedures are applied.
We have provided a detailed elaboration of the above steps (§3.1 and §3.2) and compared our method with the previous approaches (§3.3).
@@
Figure 3: Illustration of our methods. Constructing STD: We employed Prompt1 (P1) to prompt GPT-4 for generating responses, which were subsequently curated through manual screening to construct the question set for STD. Subsequently, models such as GPT-4 were engaged in a “role-playing” manner using Prompt2 (P2) to generate answers with diverse personality traits tailored to this question set, thus forming the answer set for STD; Building UPL Unsupervisedly: Initially, we employ the tokenizer of LLMs (sp(·)) to tokenize each answer text in the answer set of STD. Subsequently, we categorize the obtained sub-words based on the personality trait theme to which the answer belongs. Within each personality trait category, cool-toned words signify a low expression level of that trait, while warm-toned words indicate a high expression level. On this basis, UPL is built through two steps: iterate assignment and normalization scaling; Manipulating LLM’s personality: We employ Prompt3 (P3) to prompt the model to answer the question set of STD. During the model’s decoding process utilizing Top-p nucleus sampling, we used UPL to change the probability vector of the next prediction word. Finally, we changed the expression degree of personality traits of the model’s answers.
3.1 Building UPL Unsupervisedly
@@
When building UPL, we employ the answer set of STD, which comprises $5$ subsets with personality traits in the Big Five. For each subset, there are $200$ response texts divided into $2$ parts, that belong to high and low sub-traits respectively.
@@
We denote this answer set as $A$, the tokenizer of the target model as $sp(\cdot)$, the vocabulary of the model as $V$, and UPL as $L$. The UPL is formally presented as follows:
$L=\{L_{k}:L_{v}\}$ (1)
@@
where $L_{k}$ is initialized using $V$, and the initial values of $L_{v}$ are zero lists of length $5$, corresponding $5$ personality traits.
The construction process comprises two stages: Iterate Assignment and Normalization Scaling.
In the first stage, we tokenize the text $A_{ij}$ in $A$:
$sp(A_{ij})=\{w_{1},w_{2},\ldots\}$ (2)
where $i$ corresponds to $5$ sub-traits with $2$ polarities ($i\in\{0,1,\ldots,9\}$), and $j\in\{0,1,\ldots,199\}$.
@@
Then, we define the personality trait index $t=\frac{i}{2}$. Finally, we perform the iterate assignment on $L_{k:v}^{t}$, as follows:
$L^{t}_{[w]}=L^{t}_{[w]}+\begin{cases}+1&\text{if }i\%2=0\\ -1&\text{else}\end{cases}$ (3)
@@
$L^{t}_{v}=\{\underbrace{v_{1},v_{2},\ldots,v_{m}}_{positive},\underbrace{v_{m+1},\dots v_{m+n}}_{negative}\}$ (4)
where $m+n=|V|$. For the personality trait $t$, the averages of its positive subset and negative subset of $L^{t}_{v}$ can be expressed as follows:
$\text{Avg}^{+}(t)=\frac{1}{m}\sum_{z=1}^{|V|}\max(0,L^{t}_{v,z})$ (5)
$\text{Avg}^{-}(t)=\frac{1}{n}\sum_{z=1}^{|V|}\min(0,L^{t}_{v,z})$ (6)
The second stage is Normalization Scaling $L$. We define a hyperparameter $M$ to control the degree of normalization and scaling. A detailed ablation study on this parameter is presented in Appendix A For the personality trait $t$, this step would lead to the following inequality:
$\max\{|Avg^{-}(t)-M_{t}|,|Avg^{+}(t)-M_{t}|\}\leq\epsilon$ (7)
where $\epsilon$ has a default value of $1\times 10^{-3}$.
To achieve this inequality, we transform the value of $L$ according to Equation 8.
$L^{t}_{v}\xrightarrow{(N_{t},S_{t})}F(L^{t}_{v},N_{t},S_{t})$ (8)
where the parameter combination $(N,S)$ is obtained through the binary search algorithm. Specifically, Equation 8 is expressed as follows: